6. Using the same traits. in a cross hetneen a red, straight, ebony, shaven female and a completely unknown male. the offspriWrite your cross outline in this format Define your genotypes by matching them w/ phenoty pes X F1

6. Using the same traits, in a cross between a red, straight, ebony, shaven female, and a completely unknown male, the offspring are:

3/8 red straight gray shaven females

1/8 red curved gray shaven females

3/16 red straight gray shaven males

1/16 red curved gray shaven males

3/16 white straight gray shaven males

1/16 white curved gray shaven males

In the space below, I've started a cross outline for you. you can mark this outline up so think through the problem. and then complete your on complete outline of this cross on the next page. Include genotypes, phenotypes and ratios as appropriate to fully explain the cross ( you need to be able to draw this entire outline in an easily Readable format starting with a blank page. Only your cross outline goes on the next page, please don't write extra information or show that is not part of the cross outline).
Answer #1

Answers - Since in the offsprings all the progeny are grey is shaven this, both of these are the dominant traits. Now, No. of* Now Cat C => | C ºCc7. & le straight = 3 - 381 e curved =1. * Wwx ww 3:1 Red: white. So, now ratios of progeny :- Red Strai